{\small

\begin{thebibliography}{99}

%\bibitem[FGA]{FGA} A.\ Grothendieck: \textit{Fondements de la G{\'e}o\-m{\'e}trie Alg{\'e}\-brique}, Extraits du S{\'e}mi\-naire Bourbaki 1957--1962, Secr{\'e}tariat math{\'e}matique, Paris 1962. 


%\bibitem[EGA]{EGA} A.\ Grothendieck: \textit{{\'E}lements de G{\'e}o\-m{\'e}trie Alg{\'e}\-brique}, Publ. Math. IHES \textbf{4}, \textbf{8}, \textbf{11}, \textbf{17}, \textbf{20}, \textbf{24}, \textbf{28}, \textbf{32}, Bures-Sur-Yvette, 1960--1967; see also Grundlehren \textbf{166}, Springer-Verlag, Berlin etc.\ 1971.


%\bibitem[SGA 1]{SGA1} A.\ Grothendieck: \textit{Rev{\^e}tements {\'e}tales et groupe fondamentale}, LNM \textbf{224}, Springer-Verlag, Berlin-Heidelberg 1971. 

 
%\bibitem[SGA 3]{SGA3} M.\ Demazure, A.\ Grothendieck: \textit{Sch\'emas en Groupes I, II, III}, LNM \textbf{151}, \textbf{152}, \textbf{153}, Springer-Verlag, Berlin etc.\ 1970.


%\bibitem[SGA 7]{SGA7} P.\ Deligne, A.\ Grothendieck, et al.: \textit{S\'eminaire de G\'eom\'etrie Alg\'ebrique 7: Groupes de monodromie en g\'eom\'etrie alg\'ebrique\/}, LNM 288, Springer, Berlin-Heidelberg 1972.


%\bibitem{Anderson} G.\ Anderson: $t$-Motives, \textit{Duke Math. J.\/} \textbf{53} (1986), 457--502.


%\bibitem{Anderson2} G.\ Anderson: On Tate Modules of Formal $t$-Modules, \textit{Internat.\ Math.\ Res.\ Notices} \textbf{2} (1993), 41--52.


%\bibitem{Andreatta} F.\ Andreatta: \textit{Generalized ring of norms and generalized $(\phi,\Gamma)$-modules}, University of Padova, Preprint 2005.


%\bibitem{AB} F.\ Andreatta, O.\ Brinon: \textit{Surconvergence des r\'epresentations $p$-adiques: le cas relativ}, Preprint April 2006.


%\bibitem{Ax} J.\ Ax: Zero of polynomials over local fields -- the Galois action, \textit{J.\ Algebra} \textbf{15} (1970), 417--428.


%\bibitem{Berger02} L.\ Berger: Repr\'esentations $p$-adiques et \'equations diff\'erentielles, \textit{Invent.\ Math.\/} \textbf{148} (2002), 219--284.


\bibitem{Berger04} L.\ Berger: \textit{\'Equations diff\'erentielles $p$-adiques et $(\phi,N)$-modules filtr\'es}, Preprint on {\tt arXiv:math.NT/0406601}, version from June 29th, 2004.


\bibitem{Berger04b} L.\ Berger: Limites de repr\'esentations cristallines, \emph{Compos.\ Math.\/} {\bfseries 140} (2004), no.\ 6, 1473--1498. 


\bibitem{Berkovich} V.G.\ Berkovich: \textit{Spectral theory and analytic geometry over non-Archimedean fields}, Mathematical Surveys and Monographs \textbf{33}, Amer.\ Math.\ Soc., Providence 1990.


\bibitem{Berkovich2} V.G.\ Berkovich: {\'E}tale cohomology for non-Archimedean analytic spaces, \textit{Publ.\ Math.\ I.H.E.S.\/} \textbf{78} (1993), 5--161.


%\bibitem{Berkovich3} V.G.\ Berkovich: Smooth $p$-adic analytic spaces are locally contractible, \textit{Invent.\ Math.\/} \textbf{137} (1999), 1--84.


\bibitem{Berthelot96} P.\ Berthelot: \emph{Cohomologie rigide et cohomologie rigide \`a support propre}, prepublication 96-03, IRMA, Rennes, 1996.


%\bibitem{Blum-Stuhler} A.\ Blum, U.\ Stuhler: Drinfeld Modules and Elliptic Sheaves, in \textit{Vector Bundles on Curves -- New Directions}, pp.\ 110--188, LNM \textbf{1649}, Springer-Verlag, Berlin etc.\ 1991.


%\bibitem{BH} G.\ B{\"o}ckle, U.\ Hartl: Uniformizable Families of $t$-motives, Preprint 2004, available from {\tt arXiv:math.NT/0411262}.


%\bibitem{BoHa} M.\ Bornhofen, U.\ Hartl: \textit{Abelian Sheaves and Pure Anderson Motives over Finite Fields}, Preprint on {\tt arXiv:math.NT/0609733}.


%\bibitem{Bosch} S.\ Bosch: \textit{Lectures on Formal and Rigid Geometry}, Preprint \textbf{378}, University of M\"unster, SFB 478--Preprint Series, M\"unster 2005, available at \\ {\tt http://wwwmath1.uni-muenster.de/sfb/about/publ/heft378.ps}\,.


%\bibitem{BGR} S.\ Bosch, U.\ G{\"u}ntzer, R.\ Remmert: \textit{Non-Archimedean Analysis}, Grundlehren \textbf{261}, Springer-Verlag, Berlin etc.\ 1984.


%\bibitem{FRG} S.\ Bosch, W.\ L{\"u}tkebohmert: Formal and Rigid Geometry 
%I.\ Rigid Spaces, \textit{Math.\ Ann.\/} \textbf{295} (1993), 291--317. %\\
%II.\ Flattening Techniques, \textit{Math.\ Ann.\/} \textbf{296} (1993), 403--429. 
%III.\ The Relative Maximum Principle, \textit{Math.\ Ann.\/} \textbf{302} (1995), 1--29. \\
%IV.\ The Reduced Fiber Theorem, \textit{Invent.\ Math.\/} \textbf{119} (1995), 361--398.


%\bibitem{TG} N.\ Bourbaki: \textit{Topologie G\'en\'erale}, Hermann, Paris 1971.


%\bibitem{Bourbaki} N.\ Bourbaki: \textit{Alg\`ebre Commutative}, Hermann, Paris 1967.

 
%\bibitem{BC}  J.-F.\ Boutot, H.\ Carayol: Uniformisation $p$-adique des courbes de Shimura: les th{\'e}or{\`e}mes de \v{C}erednik et de Drinfel'd, in \textit{Courbes modulaires et courbes de Shimura (Orsay, 1987/1988)}.  Ast{\'e}risque \textbf{196-197} (1991), 45--158.


%\bibitem{Boyer00} P.\ Boyer: Mauvaise r\'eduction des vari\'et\'es de Drinfeld et correspondance de Langlands locale, \textit{Invent.\ Math.\/} \textbf{138} (1999), no.\ 3, 573--629. 


%\bibitem{Breuer} F.\ Breuer: \textit{Special subvarieties of Drinfeld modular varieties}, Preprint on {\tt arXiv:math.NT/0503452}, version from Jan.\ 11, 2007.


%\bibitem{Breuil1} C.\ Breuil: \textit{Schemas en groupes et corps des normes}, unpublished (1998), 13 pages, available at {\tt http://www.ihes.fr/$\sim$breuil/}.


\bibitem{Breuil2} C.\ Breuil: Groupes $p$-divisibles, groupes finis et modules filtr\'es, \textit{Ann.\ of Math.\/} \textbf{152} (2000), 489--549.


%\bibitem{Brinon} O.\ Brinon: \textit{Repr\'esentations Galoisiennes $p$-adiques dans le cas relative}, Th\`ese de l'Universit\'e de Paris 11 (Orsay), 2004.


%\bibitem{Cerednik} V.\ \v{C}erednik: Uniformization of algebraic curves by discrete arithmetic subgroups of $\PGL_2(k_w)$ with compact quotients,  \textit{Math.\ USSR.\ Sb.\/} \textbf{29} (1976), 55--78.


%\bibitem{CC} F.\ Cherbonnier, P.\ Colmez: Repr\'esentations $p$-adiques surconvergentes, \textit{Invent.\ Math.\/} \textbf{133} (1998), no.\ 3, 581--611.


%\bibitem{CC2} F.\ Cherbonnier, P.\ Colmez: Th\'eorie d'Iwasawa des repr\'esentations $p$-adiques d'un corps local, \textit{J.\ Amer.\ Math.\ Soc.\/} \textbf{12} (1999), 241--268.


%\bibitem{Colmez1} P.\ Colmez: Espaces de Banach de dimension finie, \textit{J.\ Inst.\ Math.\  Jussieu} \textbf{1} (2002), 331--439.


%\bibitem{Colmez02} P.\ Colmez: Espaces de Banach de dimension finie, \textit{J.\ Inst.\ Math.\ Jussieu} \textbf{1} (2002), no.\ 3, 331--439.


\bibitem{Colmez} P.\ Colmez: \textit{Espaces Vectoriels de dimension finie et r\'epresentations de de Rham}, Preprint 2003, available at {\tt http://www.math.jussieu.fr/$\sim$colmez}.


\bibitem{CF} P.\ Colmez, J.-M.\ Fontaine: Construction des repr\'esentations $p$-adiques semi-stables, \textit{Invent.\ Math.\/} \textbf{140} (2000), no. 1, 1--43.


%\bibitem{FLT} G.\ Cornell, J.\ Silverman, G.\ Stevens, eds.: \textit{Modular forms and Fermat's last theorem}, Springer-Verlag, New York, 1997.


%\bibitem{DH} P.\ Deligne, D.\ Husem{\"o}ller: Survey of Drinfeld modules, in \textit{Current trends in arithmetical algebraic geometry (Arcata, Calif., 1985)}, Contemp.\ Math.\ \textbf{67}, Amer.\ Math.\ Soc., Providence 1987; pp.\ 25--91.


%\bibitem{DI} P.\ Deligne, L.\ Illusie: Rel\`evements modulo $p^2$ et d\'ecomposition du complexe de de Rham, \textit{Invent.\ Math.\/} \textbf{89} (1987), 247--270.


%\bibitem{DMilne} P.\ Deligne, J.\ Milne: Tannakian Categories, in \textit{Hodge Cycles, Motives, and Shimura Varieties}, pp.\ 101--228, LNM \textbf{900}, Springer-Verlag, New York 1982.


%\bibitem{DM} P.\ Deligne, D.\ Mumford: The Irreducibility of the Space of Curves of Given Genus, \textit{Publ.\ Math.\ IHES\/} \textbf{36} (1969), 75--110.


%\bibitem{Demazure} M.\ Demazure: \textit{$p$-divisible groups}, LNM \textbf{302}, Springer-Verlag, Berlin-New York, 1972.


%\bibitem{Drinfeld} V.G.\ Drinfeld: Elliptic Modules, \textit{Math.\ USSR-Sb.\/} \textbf{23} (1976), 561--592.


\bibitem{Drinfeld2} V.G.\ Drinfeld: Coverings of $p$-adic symmetric domains, \textit{Funct.\ Anal.\ Appl.\/} \textbf{10} (1976), 107--115.


%\bibitem{Drinfeld3} V.G.\ Drinfeld: Commutative Subrings of Certain Noncommutative Rings, \textit{Funct.\ Anal.\ Appl.\/} \textbf{11} (1977), 9--12.


%\bibitem{Drinfeld4} V.G.\ Drinfeld: A proof of Langlands' global conjecture for ${\rm GL}(2)$ over a function field, \textit{Funct.\ Anal.\ Appl.\/} \textbf{11} (1977), no. 3, 223--225.


%\bibitem{Drinfeld5} V.G.\ Drinfeld: Moduli variety of $F$-sheaves, \textit{Funct.\ Anal.\ Appl.\/} \textbf{21} (1987), no. 2, 107--122.


%\bibitem{Eisenbud} D.\ Eisenbud: \textit{Commutative Algebra with a View Toward Algebraic Geometry}, GTM \textbf{150}, Springer-Verlag, Berlin etc.\ 1995.


%\bibitem{Faltings} G.\ Faltings: Endlichkeitss{\"a}tze f{\"u}r abelsche Variet{\"a}ten {\"u}ber Zahlk{\"o}rpern, \textit{Invent.\ Math.\/} \textbf{73} (1983), 349--366.


%\bibitem{Faltings1} G.\ Faltings: $p$-adic Hodge theory, \textit{J.\ Amer.\ Math.\ Soc.\/} \textbf{1}  (1988),  no.\ 1, 255--299.


%\bibitem{Faltings} G.\ Faltings: Crystalline cohomology and $p$-adic Galois-representations, \textit{Algebraic analysis, geometry, and number theory (Baltimore 1988)}, pp.\ 25--80, Johns Hopkins Univ.\ Press, Baltimore 1989.


%\bibitem{Faltings99} G.\ Faltings: Integral crystalline cohomology over very ramified valuation rings, \textit{J.\ Amer.\ Math.\ Soc.\/} \textbf{12} (1999), no.\ 1, 117--144.


%\bibitem{Faltings02} G.\ Faltings: Almost \'etale extensions, \textit{Cohomologies $p$-adiques et applications arithm\'etiques, II\/}, Ast\'erisque \textbf{279} (2002), 185--270. 


%\bibitem{Faltings95} G.\ Faltings: Mumford-Stabilit\"at in der algebraischen Geometrie, \textit{Proceedings of the International Congress of Mathematicians, Vol.\ 1, 2 (Zürich, 1994)\/}, pp.\ 648--655, Birkh\"auser, Basel, 1995. 


\bibitem{Faltings07} G.\ Faltings: \textit{Coverings of $P$-adic Period Domains}, Preprint 2007.


%\bibitem{Fargues04} L.\ Fargues: Cohomologie des espaces de modules de groupes $p$-divisibles et correspondances de Langlands locales, \textit{Vari\'et\'es de Shimura, espaces de Rapoport-Zink et correspondances de Langlands locales,  Ast\'erisque} \textbf{291} (2004), 1--199. 


\bibitem{Fontaine79} J.-M.\ Fontaine: Modules galoisiens, modules filtr\'es et anneaux de Barsotti-Tate, in \textit{Journ\'ees de g\'eometrie algebrique de Rennes (III), Ast\'erisque} \textbf{65} (1979), 3--80.


%\bibitem{Fontaine82} J.-M.\ Fontaine: Sur certains types de repr\'esentations $p$-adiques du groupe de Galois d'un corps local; construction d'un anneau de Barsotti-Tate, \textit{Ann.\ of Math.\ (2)} \textbf{115}  (1982), no.\ 3, 529--577.


%\bibitem{Fontaine94} J.-M.\ Fontaine: Le corps des p\'eriodes $p$-adiques, \textit{P\'eriodes $p$-adiques (Bures-sur-Yvette, 1988),  Ast\'erisque} \textbf{223} (1994), 59--111.


%\bibitem{Fontaine94b} J.-M.\ Fontaine: Repr\'esentations $p$-adiques semi-stables, \textit{P\'eriodes $p$-adiques (Bures-sur-Yvette 1988)\/}, pp.\ 113--184, Ast\'erisque \textbf{223}, Soc.\ Math.\ France, Paris 1994.


%\bibitem{FM} J.-M.\ Fontaine, W.\ Messing: $p$-adic periods and $p$-adic étale cohomology, \textit{Current trends in arithmetical algebraic geometry (Arcata, Calif., 1985)\/}, pp.\ 179--207, Contemp.\ Math.\ 67, Amer.\ Math.\ Soc., Providence, RI, 1987. 


%\bibitem[FR]{Fontaine-Rapoport} J.-M.\ Fontaine, M.\ Rapoport: Existence de filtrations admissibles sur des isocristaux, \textit{Bull.\ Soc.\ Math.\ France\/} \textbf{133} (2005), no.\ 1, 73--86. 


%\bibitem{FW1} J.-M.\ Fontaine, J.-P.\ Wintenberger:  Le ``corps des normes'' de certaines extensions alg\'ebriques de corps locaux, \textit{C.\ R.\ Acad.\ Sci.\ Paris S\'er.\ A-B} \textbf{288} (1979), no.\ 6, 367--370.


%\bibitem{FW2} J.-M.\ Fontaine, J.-P.\ Wintenberger: Extensions alg\'ebrique et corps des normes des extensions APF des corps locaux, \textit{C.\ R.\ Acad.\ Sci.\ Paris S\'er.\ A-B} \textbf{288} (1979), no.\ 8, 441--444.


%\bibitem{Fresnel} J.\ Fresnel, M.\ van der Put: \textit{G{\'e}om{\'e}trie analytique rigide et applications}, Progress in Mathematics \textbf{218}, Birkh{\"a}user, Basel 2004.


%\bibitem{GardeynDiss} F.\ Gardeyn: \textit{$t$-Motives and Galois Representations}, Dissertation Universiteit Gent, Oct.\ 2001.


%\bibitem{Gardeyn2} F.\ Gardeyn: A Galois criterion for good reduction of $\tau$-sheaves, \textit{J.\ Number Theory\/} \textbf{97} (2002), 447--471.


%\bibitem{Gardeyn3} F.\ Gardeyn: New criteria for uniformization of t -motives. Preprint 2001.


%\bibitem{Gardeyn4} F.\ Gardeyn: The structure of analytic $\tau$-sheaves, \textit{J.\ Number Th.\/} \textbf{100} (2003), 332--362.


%\bibitem{Genestier} A.\ Genestier: \textit{Espaces sym{\'e}triques de Drinfeld\/}, Ast{\'e}risque \textbf{234}, Soc.\ Math.\ France, Paris 1996.


%\bibitem{GL} A.\ Genestier, V.\ Lafforgue: \textit{Th\'eories de Fontaine et Fontaine-Laffaille en \'egale charact\'eristique\/}, Preprint 2004.


%\bibitem{Gerritzen} L.\ Gerritzen: Zerlegungen der Picard-Gruppe nichtarchimedischer holomorpher R\"aume, \textit{Comp.\ Math.\/} \textbf{35} no.\ 1 (1977), 23--38.


%\bibitem{Goss} D.\ Goss: \textit{Basic Structures of Function Field Arithmetic}, Ergebnisse \textbf{35}, Springer-Verlag, Berlin-Heidelberg-New York 1996.


%\bibitem{GossDict} D.\ Goss: Dictionary, \textit{The arithmetic of function fields (Columbus, OH, 1991)},  pp.\ 475--482, Ohio State Univ.\ Math.\ Res.\ Inst.\ Publ., 2, de Gruyter, Berlin, 1992.


\bibitem{Grothendieck2} A.\ Grothendieck: Groupes de Barsotti-Tate et cristaux,  \textit{Actes du Congr\`es International des Math\'ematiciens (Nice, 1970), Tome 1},  pp.\ 431--436, Gauthier-Villars, Paris, 1971. 


%\bibitem{GrothendieckCristaux} A.\ Grothendieck: Groupes de Barsotti-Tate et cristaux de Dieudonn{\'e}, \textit{S{\'e}minaire de Math{\'e}matiques Sup{\'e}rieures} \textbf{45}, Les Presses de l'Universit{\'e} de Montr{\'e}al, Montreal 1974. 


%\bibitem{Harris} M.\ Harris, R.\ Taylor: \textit{The geometry and cohomology of some simple Shimura varieties}, Ann.\ Math.\ Stud. \textbf{151}, Princeton University Press, Princeton, 2001.


%\bibitem{Harder} G.\ Harder: Wittvektoren, \textit{Jahresber.\ Deutsch.\ Math.-Verein.\/} \textbf{99} (1997), no.\ 1, 18--48.


%\bibitem{HartlAbSh} U.\ Hartl: Uniformizing the Stacks of Abelian Sheaves, in \textit{Number Fields and Function fields - Two Parallel Worlds, Papers from the 4th Conference held on Texel Island, April 2004}, G.\ van der Geer, B.\ Moonen, R.\ Schoof, Editors, pp.\ 167--222, Progress in Math.\ 239, Birkh\"auser-Verlag, Basel 2005. See also {\tt arXiv:math.NT/0409341}. 


\bibitem{HartlPSp} U.\ Hartl: \textit{Period Spaces for Hodge Structures in Equal Characteristic}, Preprint on {\tt arXiv:math.NT/0511686}.


\bibitem{HartlRZ} U.\ Hartl: \textit{On a Conjecture of Rapoport and Zink}, Preprint on {\tt arXiv:math.NT/0605254}.


\bibitem{HartlDict} U.\ Hartl: \textit{A Dictionary between Fontaine-Theory and its Analogue in Equal Characteristic}, Preprint on {\tt arXiv:math.NT/0607182}.


%\bibitem{Crystals} U.\ Hartl: \textit{Local Shtuka and Divisible Local Anderson Modules}, in preparation.


%\bibitem{FormalStacks} U.\ Hartl: Formal Algebraic Stacks, in preparation.


%\bibitem{HP} U.\ Hartl, R.\ Pink: Vector bundles with a Frobenius structure on the punctured unit disc, \textit{Comp.\ Math.\/} \textbf{140} n.3 (2004), 689--716. 


%\bibitem{Hartshorne} R.\ Hartshorne: \textit{Algebraic Geometry}, GTM \textbf{52}, Springer-Verlag, Berlin etc.\ 1977.


%\bibitem{Hausberger} T.\ Hausberger: Uniformisations des Vari{\'e}t{\'e}s de Laumon-Rapoport-Stuhler et conjecture de Drinfeld-Carayol, \textit{Ann.\ Inst.\ Fourier (Grenoble)} \textbf{55} (2005), no.\ 4, 1285--1371.



%\bibitem{HG1} M.J.\ Hopkins, B.H.\ Gross: The rigid analytic period mapping, Lubin-Tate space, and stable homotopy theory, \textit{Bull.\ Amer.\ Math.\ Soc.\/} \textbf{30} (1994), no. 1, 76--86.

 
%\bibitem{HG2} M.J.\ Hopkins, B.H.\ Gross: Equivariant vector bundles on the Lubin-Tate moduli space, in \textit{Topology and representation theory (Evanston, IL, 1992)}, pp.\ 23--88, Contemp.\ Math.\ \textbf{158}, Amer.\ Math.\ Soc., Providence, RI, 1994.


%\bibitem{Illusie} L.\ Illusie: Cohomologie de de Rham et cohomologie \'etale $p$-adique (d'apr\`es G.\ Faltings, J.-M.\ Fontaine et al.), \textit{S\'eminaire Bourbaki, Vol. 1989/90,  Ast\'erisque} \textbf{189-190} (1990), Exp.\ No.\ 726, 325--374.


\bibitem{dJ} J.\ de Jong: {\'E}tale Fundamental groups of non-Archimedean analytic spaces, \textit{Comp.\ Math.\/} \textbf{97} (1995), 89--118.


%\bibitem{dJ95} J.\ de Jong: Crystalline Dieudonn\'e module theory via formal and rigid geometry, \textit{Inst.\ Hautes \'Etudes Sci.\ Publ.\ Math.\/} \textbf{82} (1995), 5--96.


%\bibitem{dJ98} A.J.\ de Jong: Homomorphisms of Barsotti-Tate groups and crystals in positive characteristic, \textit{Invent.\ Math.\/} \textbf{134} (1998), no.\ 2, 301--333. 


%\bibitem{JongPut} J.\ de Jong, M.\ van der Put: {\'E}tale Cohomology of Rigid Analytic Spaces, \textit{Doc.\ Math.\/} \textbf{1} (1996), 1--56.


%\bibitem{Katz1} N.\ Katz: $p$-adic properties of modular schemes and modular forms, \textit{Modular functions of one variable, III (Proc.\ Internat.\ Summer School, Univ.\ Antwerp, Antwerp, 1972)},  pp.\ 69--190. LNM \textbf{350}, Springer, Berlin, 1973. 


%\bibitem{Katz} N.\ Katz: Slope filtration of $F$-crystals, \textit{Journ{\'e}es de G{\'e}om{\'e}trie Alg{\'e}brique de Rennes (Rennes, 1978), Vol. I},  pp. 113--163, Ast{\'e}risque \textbf{63}, Soc.\ Math.\ France, Paris 1979.


%\bibitem{Katz-Mazur} N.\ Katz, B.\ Mazur: \textit{Arithmetic Moduli of Elliptic Curves}, Ann.\ Math.\ Stud. \textbf{108}, Princeton University Press, Princeton 1985. 

 
%\bibitem{Kedlaya2} K.\ Kedlaya: A $p$-adic local monodromy theorem, \textit{Ann.\ of Math.\ (1)} \textbf{160} (2004), 93--184.


\bibitem{Kedlaya} K.\ Kedlaya:  Slope filtrations revisited, \textit{Doc.\ Math.\/} \textbf{10} (2005), 447--525.


%\bibitem{KiehlEndl} R.\ Kiehl: Der Endlichkeitssatz f{\"u}r eigentliche Abbildungen in der nichtarchimedischen Funktionentheorie, \textit{Invent.\ Math.\/} \textbf{2} (1967), 191--214.


%\bibitem{KiehlAB} R.\ Kiehl: Theorem A und B in der nichtarchimedischen Funktionentheorie, \textit{Invent.\ Math.\/} \textbf{2} (1967), 256--273.


\bibitem{Kisin06} M.\ Kisin: Crystalline Representations and $F$-Crystals, \textit{Algebraic geometry and number theory}, pp.\ 459--496, Progr.\ Math., 253, Birkh\"auser Boston, Boston, MA, 2006. 


%\bibitem{Knutson} D.\ Knutson: \textit{Algebraic Spaces}, LNM \textbf{203}, Springer-Verlag, Berlin etc.\ 1971.


%\bibitem{Kottwitz85} R.\ Kottwitz: Isocrystals with additional structure, \textit{Compositio Math.\/} \textbf{56}  (1985),  no. 2, 201--220.


%\bibitem[Ko2]{Kottwitz97} R.\ Kottwitz: Isocrystals with additional structure II, \textit{Compositio Math.\/} \textbf{109} (1997), no.\ 3, 255--339.


%\bibitem{Lafforgue2} L.\ Lafforgue: \textit{Chtoucas de Drinfeld et conjecture de Ramanujan-Petersson}, Ast{\'e}risque \textbf{243}, Soc.\ Math.\ France, Paris 1997.


%\bibitem{Lafforgue1} L.\ Lafforgue: Une compactification des champs classifiant les chtoucas de Drinfeld, \textit{J.\ Amer.\ Math.\ Soc.\/} \textbf{11}  (1998),  no.\ 4, 1001--1036.


%\bibitem{Lafforgue} L.\ Lafforgue: Chtoucas de Drinfeld et correspondance de Langlands, \textit{Invent.\ Math.\/} \textbf{147} (2002), 1--241.


%\bibitem{Langton} S.\ Langton: Valuative criteria for families of vector bundles on algebraic varieties, \textit{Ann.\ of Math.\ (2)\/} \textbf{101} (1975), 88--110.


%\bibitem{Laumon} G.\ Laumon: \textit{Cohomology of Drinfeld Modular Varieties I\/}, Cambridge Studies in Advanced Mathematics \textbf{41}, Cambridge University Press, Cambridge, 1996.


%\bibitem{Laumon97} G.\ Laumon: Drinfeld shtukas, in \textit{Vector bundles on curves---new directions (Cetraro, 1995)}, pp.\ 50--109, LNM \textbf{1649}, Springer, Berlin, 1997.


%\bibitem{LaumonMB} G.\ Laumon, L.\ Moret-Bailly: \textit{Champs alg{\'e}briques\/}, Ergebnisse \textbf{39}, Springer-Verlag, Berlin etc.\ 2000.


%\bibitem{LRS} G.\ Laumon, M.\ Rapoport, U.\ Stuhler: ${\cal D}$-elliptic sheaves and the Langlands correspondence, \textit{Invent.\ Math.\/} \textbf{113} (1993), 217--338.


%\bibitem{Lazard} M.\ Lazard: Les z\'eros des fonctions analytiques d'une variable sur un corps valu\'e complet, \textit{Inst. Hautes \'Etudes Sci.\ Publ.\ Math.\/} \textbf{14} (1962), 47--75.


%\bibitem{LT} J.\ Lubin, J.\ Tate: Formal complex multiplication in local fields, \textit{Ann.\ of Math.\/} (2) \textbf{81} (1965), 380--387.


%\bibitem{Lubo} W.\ L{\"u}tkebohmert: Formal-algebraic and rigid-analytic geometry, \textit{Math.\ Ann.\/} \textbf{286} (1990), 341--371.


%\bibitem{Lubo1} W.\ L{\"u}tkebohmert: Vektorraumb{\"u}ndel {\"u}ber nichtarchimedischen holomorphen R{\"a}umen, \textit{Math.\ Z.\/} \textbf{152} (1977), 127--143.


%\bibitem{Lubo2} W.\ L{\"u}tkebohmert: Riemann's existence problem for a $p$-adic field, \textit{Invent.\ Math.\/} \textbf{111} (1993), 309 \- 330.


%\bibitem{Manin} Y.I.\ Manin: The theory of commutative formal groups over fields of finite characteristic, \textit{Russian Math. Surveys} \textbf{18} No.~6 (1963), 1--83.


%\bibitem{Matzat} H.\ Matzat: Introduction to Drinfeld modules, in \textit{Drinfeld modules, modular schemes and applications (Alden-Biesen, 1996)}, pp.\ 3--16, World Sci.\ Publishing, River Edge, NJ, 1997. 


\bibitem{Messing} W.\ Messing: \textit{The Crystals Associated to Barsotti-Tate Groups}, LNM \textbf{264}, Springer-Verlag, Berlin etc.\ 1972.


%\bibitem{GIT} D.\ Mumford, J.\ Fogarty, F.\ Kirwan: \textit{Geometric invariant theory}, Ergebnisse \textbf{34}, Springer-Verlag, Berlin, 1994.


%\bibitem[Ni]{Niziol98} W.\ Niziol: Crystalline conjecture via $K$-theory, \textit{Ann.\ Sci.\ \'Ecole Norm.\ Sup.\ (4)} \textbf{31} (1998), no.\ 5, 659--681.


%\bibitem[Or1]{Orlik05} S.\ Orlik: The cohomology of period domains for reductive groups over local fields, \textit{Invent.\ Math.\/} \textbf{162} (2005), no.\ 3, 523--549.


%\bibitem[Or2]{Orlik05b} S. Orlik: \textit{The continuous cohomology of period domains over local fields}, Preprint 2005 on {\tt arXiv:math.NT/0505409}.


%\bibitem{Pink} R.\ Pink: \textit{Hodge Structures over Function Fields}, Preprint 1997, available at {\tt http://www.math.ethz.ch/$\sim$pink}.


%\bibitem{Potemine} I.Y.\ Potemine: Drinfeld-Anderson Motives and Multicomponent KP Hierarchy, in \textit{Recent progress in algebra (Taejon/Seoul 1997)},  Contemp.\ Math.\ \textbf{224}, Amer.\ Math.\ Soc., Providence 1999.


%\bibitem{Put} M.\ van der Put: Cohomology on affinoid spaces, \textit{Comp.\ Math.\/} \textbf{45} (1982), 165--198.


%\bibitem{PutSchneider} M.\ van der Put, P. Schneider: Points and topologies in rigid geometry, \textit{Math.\ Ann.\/} \textbf{302} (1995), 81--103.


%\bibitem{Quillen} D.\ Quillen: Higher algebraic K-theory I, in \textit{Algebraic K-Theory, I: Higher $K$-theories (Proc.\ Conf., Battelle Memorial Inst., Seattle, Wash., 1972)},  pp. 85--147, LNM \textbf{341}, Springer-Verlag, Berlin 1973.


%\bibitem[Ra]{Rapoport95} M.\ Rapoport: Non-Archimedean period domains, \textit{Proceedings of the International Congress of Mathematicians, Vol.\ 1 (Zürich, 1994)}, pp.\ 423--434, Birkh\"auser, Basel, 1995.


\bibitem{RZ} M.\ Rapoport, T.\ Zink: \textit{Period Spaces for $p$-divisible Groups\/}, Ann.\ Math.\ Stud.\ \textbf{141}, Princeton University Press, Princeton 1996.


%\bibitem{Raynaud} M.\ Raynaud: G{\'e}om{\'e}trie analytique rigide d'apres Tate, Kiehl, \ldots, \textit{Bul.\ Soc.\ Math.\ France, M{\'e}moires \/} \textbf{39/40} (1974), 319--327.


%\bibitem{Rosen} M.\ Rosen: Formal Drinfeld modules, \textit{J.\ Number Theory\/} \textbf{103}  (2003), 234--256.


%\bibitem{Rotman} J.\ Rotman: \textit{An Introduction to Homological Algebra}, Pure and Applied Mathematics \textbf{85}, Academic Press, New York, 1979.


%\bibitem{Saidi} M.\ Saidi, \textit{Moduli schemes of Drinfeld modules} in ``Drinfeld Modules, Modular Schemes and Applications'', edited by E.-U.\ Gekeler et al., World Scientific, 1997, 17--31.


%\bibitem{Schneider} P.\ Schneider: Points of rigid analytic varieties, \textit{J.\ reine Angew.\ Math.\/} \textbf{434} (1993), 127--157.


%\bibitem{SchnStuhler} P.\ Schneider, U.\ Stuhler: The cohomology of $p$-adic symmetric spaces, \textit{Invent.\ Math.\/} \textbf{105} (1991), 47--122.


%\bibitem{Serre2} J.-P.\ Serre: Groupes alg\'ebriques associ\'es aux modules de Hodge-Tate, \textit{Journ\'ees de G\'eom\'etrie Alg\'ebrique de Rennes (Rennes, 1978)}, Vol. III, pp. 155--188, Ast\'erisque \textbf{65}, Soc.\ Math.\ France, Paris 1979. 


%\bibitem{SerreLF} J.-P.\ Serre: \textit{Local fields}, GTM \textbf{67}, Springer-Verlag, New York-Berlin, 1979.


%\bibitem{Serre} J.-P.\ Serre: \textit{Algebraic Groups and Class fields}, GTM \textbf{117}, Springer-Verlag, Berlin etc.\ 1988.


%\bibitem{Seshadri} C.S.\ Seshadri: \textit{Fibr{\'e}s vectoriels sur les courbes alg{\'e}briques\/}, Ast{\'e}risque \textbf{96}, Soc.\ Math.\ France, Paris 1982.


%\bibitem{Stuhler} U.\ Stuhler: $P$-Adic Homogeneous Spaces and Moduli Problems, \textit{Math.\ Z.\/} \textbf{192} (1986), 491--540.


%\bibitem{Taguchi93} Y.\ Taguchi: Semi-simplicity of the Galois Representations Attached to Drinfeld Modules over Fields of ``Infinite Characteristics'', \textit{J.\ Number Theory\/} \textbf{44} (1993), no.\ 3, 292--314.


%\bibitem{Taguchi95} Y. Taguchi: A duality for finite $t$-modules, \textit{J.\ Math. Sci. Univ. Tokyo} \textbf{2} (1995), 563--588.


%\bibitem{TW} Y.\ Taguchi, D.\ Wan: $L$-functions of $\phi$-sheaves and Drinfeld modules, \textit{J.\ Amer.\ Math.\ Soc.\/} \textbf{9} (1996),  no.\ 3, 755--781.


%\bibitem{Tate66} J. Tate: $p$-Divisible Groups, \textit{Proc.\ Conf.\ Local Fields (Driebergen, 1966)},  pp. 158--183, Springer-Verlag, Berlin 1967.


%\bibitem{Tate71} J.\ Tate: Rigid analytic spaces, \textit{Invent.\ Math.\/} \textbf{12} (1971), 257--289.


%\bibitem{Totaro96} B.\ Totaro: Tensor products in $p$-adic Hodge theory, \textit{Duke Math.\ J.\/} \textbf{83} (1996), no.\ 1, 79--104.


%\bibitem{Tsuji} T.\ Tsuji: $p$-adic \'etale cohomology and crystalline cohomology in the semi-stable reduction case, \textit{Invent.\ Math.\/} \textbf{137}  (1999),  no.\ 2, 233--411.


%\bibitem{Wintenberger} J.-P.\ Wintenberger: Le corps des normes de certaines extensions infinies de corps locaux; applications, \textit{Ann.\ Sci.\ \'Ecole Norm.\ Sup.\ (4)} \textbf{16}  (1983),  no.\ 1, 59--89.


%\bibitem{Wintenberger94} J.-P.\ Wintenberger: Th\'eor\`eme de comparaison $p$-adique pour les sch\'emas ab\'eliens, I. Construction de l'accouplement de p\'eriodes, \textit{P\'eriodes $p$-adiques (Bures-sur-Yvette, 1988).  Ast\'erisque} \textbf{223} (1994), 349--397.


%\bibitem{Zink} T. Zink: \textit{Cartiertheorie kommutativer formaler Gruppen}, Teubner-Texte zur Mathematik \textbf{68}, Teubner Verlagsgesellschaft, Leipzig 1984.


%\bibitem{Zink02} T.\ Zink: The display of a formal $p$-divisible group, \textit{Cohomologies $p$-adiques et applications arithm\'etiques, Vol.\ I\/}, pp.\ 127--248, Ast\'erisque \textbf{278}, Soc.\ Math.\ France, Paris 2002.
\end{thebibliography}

}

\documentclass[11pt]{article}

\usepackage{a4,amsfonts,exscale}

\long\def\forget#1{}

\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amscd}
\usepackage{textcomp}

\usepackage{amsthm}
\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{proposition}[theorem]{Proposition}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}

\newtheorem{bigexample}[theorem]{Example}

\theoremstyle{remark}

\newtheorem{remark}[theorem]{Remark}

\usepackage{xy}
\xyoption{all}

\newdir^{ (}{{}*!/-3pt/\dir^{(}}    
\newdir^{  }{{}*!/-3pt/\dir^{}}    
\newdir_{ (}{{}*!/-3pt/\dir_{(}}    
\newdir_{  }{{}*!/-3pt/\dir_{}}

\newcounter{zahl}

\def\theenumi{(\alph{enumi})}

\def\p@enumii{\theenumi}

\newcommand{\DS}{\displaystyle}
\newcommand{\TS}{\textstyle}

\newcommand{\SSC}{\scriptscriptstyle}

\DeclareMathOperator{\Frob}{Frob}

\DeclareMathOperator{\Grass}{Grass}

\DeclareMathOperator{\Hom}{Hom}

\DeclareMathOperator{\Lie}{Lie}

\DeclareMathOperator{\Var}{V}

\newcommand{\alg}{{\rm alg}}
\newcommand{\an}{{\rm an}}

\newcommand{\cris}{{\rm cris}}

\DeclareMathOperator{\id}{\,id}

\DeclareMathOperator{\ord}{ord}

\newcommand{\rig}{{\rm rig}}
\DeclareMathOperator{\rk}{rk}

\renewcommand{\phi}{\varphi}
\renewcommand{\epsilon}{\varepsilon}

\usepackage{amsfonts}
\newcommand{\BOne} {{\mathchoice{\hbox{\rm1\kern-2.7pt l\kern.9pt}}
                              {\hbox{\rm1\kern-2.7pt l\kern.9pt}}
                              {\hbox{\scriptsize\rm1\kern-2.3pt l\kern.4pt}}
                              {\hbox{\scriptsize\rm1\kern-2.4pt l\kern.5pt}}}}

\newcommand{\BD}{{\mathbb{D}}}

\newcommand{\BF}{{\mathbb{F}}}

\newcommand{\BN}{{\mathbb{N}}}

\newcommand{\BQ}{{\mathbb{Q}}}

\newcommand{\BX}{{\mathbb{X}}}

\newcommand{\BZ}{{\mathbb{Z}}}

\newcommand{\bA}{{\mathbf{A}}}
\newcommand{\bB}{{\mathbf{B}}}

\newcommand{\bE}{{\mathbf{E}}}

\newcommand{\calD}{{\mathcal{D}}}

\newcommand{\CF}{{\mathcal{F}}}
\newcommand{\CG}{{\mathcal{G}}}

\newcommand{\CM}{{\mathcal{M}}}
\newcommand{\CN}{{\mathcal{N}}}
\newcommand{\CO}{{\mathcal{O}}}

\newcommand{\CS}{{\mathcal{S}}}

\newcommand{\rbij}{\mbox{\mathsurround=0pt \;$-\hspace{-0.75em}\stackrel{\sim\quad}{\longrightarrow}$\;}}

\let\setminus\smallsetminus

\newcommand{\es}{\enspace}

\newcommand{\dual}{^{\SSC\lor}}

\newcommand{\wt}[1]{{\widetilde{#1}}}

\usepackage{ifthen}

\newcommand{\invlim}[1][]{\ifthenelse{\equal{#1}{}}
{\DS \lim_{\longleftarrow}}
{\DS \lim_{\underset{#1}{\longleftarrow}}}
}

\newcommand{\dirlim}[1][]{\ifthenelse{\equal{#1}{}}
{\DS \lim_{\longrightarrow}}
{\DS \lim_{\underset{#1}{\longrightarrow}}}
}

\newcommand{\dotBD}{\vbox{\hbox{\kern2pt\bf.}\vskip-4.5pt\hbox{$\BD$}}}

\def\?{\ 
???\ \immediate\write16{}
\immediate\write16{Warning: There was still a question mark . . . }
\immediate\write16{}}

\DeclareMathOperator{\Nilp}{\CN \!{\it ilp}}
\DeclareMathOperator{\Sets}{\CS \!{\it ets}}

\def\isoto{\rbij}

\newbox\mybox
\def\arrover#1{\mathrel{
       \setbox\mybox=\hbox spread 1.4em{\hfil$\scriptstyle#1$\hfil}
       \vbox{\offinterlineskip\copy\mybox
             \hbox to\wd\mybox{\rightarrowfill}}}}

\usepackage{amsthm}
\theoremstyle{plain}
\newtheorem{problem}[theorem]{Problem}

\begin{document}

\author{Urs Hartl
\footnote{The author acknowledges support of the Deutsche Forschungsgemeinschaft in form of DFG-grant HA3006/2-1}
}

\title{On period spaces for $p$-divisible groups}

\maketitle

\begin{abstract}
\noindent
In their book 
Rapoport and Zink constructed rigid analytic period spaces for Fontaine's filtered isocrystals, and period morphisms from moduli spaces of $p$-divisible groups to some of these period spaces. We determine the image of these period morphisms, thereby contributing to a question of Grothendieck. We give examples showing that only in rare cases the image is all of the Rapoport-Zink period space.

\medskip

\begin{center}
\bfseries R\'esum\'e
\end{center}

\noindent 
Dans leur livre, Rapoport et Zink ont construits des espaces des p\'eriodes, rigides ana\-lytiques pour les isocristaux filtr\'es de Fontaine. Egalement ils ont construits des morphismes des p\'eriodes entre des espaces modulaires des groupes de Barsotti-Tate et certaines de leurs espaces des p\'eriodes. Dans cet article nous d\'eterminons l'image des morphismes des p\'eriodes, contribuant ainsi \`a une question de Grothendieck. Nous pr\'esentons des examples montrant que l'image ne co\"incide que rarement avec tout l'espace des p\'eriodes de Rapoport-Zink.

\medskip

\noindent
{\it Mathematics Subject Classification (2000)\/}: 
11S20,  
(14G22,  
14L05,  
14M15)  
\end{abstract}

\section{A question of Grothendieck}

 Fix a Barsotti-Tate group $\BX$ over $\BF_p^\alg$ of height $h$ and dimension $d$. Let $W:=W(\BF_p^\alg)$ be the ring of Witt vectors and let $K_0:=W[\frac{1}{p}]$. Let $\CO_K$ be a complete discrete valuation ring with residue field $\BF_p^\alg$ and fraction field $K$ of characteristic $0$. To every Barsotti-Tate group $X$ over $\CO_K$ with $\BX\cong X\otimes_{\CO_K}\BF_p^\alg=:X_{\BF_p^\alg}$ the theory of Grothendieck-Messing~\cite{Messing} associates an extension
\[
\xymatrix {
0\ar[r] & (\Lie X\dual)_K\dual \ar[r] & \BD(\BX)_K\ar[r] & \Lie X_K\ar[r] & 0
}
\]
where $\BD(\BX)_K$ is the crystal of Grothendieck-Messing evaluated on $K$. We denote by $\CF$  the Grassmannian of $(h-d)$-dimensional subspaces of $\BD(\BX)_{K_0}$.

\begin{problem} \label{ProblemGroth}
(A.~Grothendieck \cite[Remarque 3, p.\ 435]{Grothendieck2})
Describe the subset of $\CF$ formed by the points $(\Lie X\dual)_K\dual$ where $X$ is any deformation of $\BX$ over any complete discrete valuation ring $\CO_K$ with residue field $\BF_p^\alg$ and fraction field $K$ of characteristic $0$.
\end{problem}

This problem is yet unsolved. However, a contribution was made by Rapoport and Zink~\cite{RZ} in the following way. Set $D:=\BD(\BX)_{K_0}$ and $\phi_D:=\BD(\Frob_\BX)_{K_0}$. Let $L_K\in\CF$ be a point defined over a complete, rank one valued extension $K$ of $K_0$ (which not necessarily is discrete). View $L_K$ as a $K$-subspace of $D_K:=D\otimes_{K_0}K$. One defines the \emph{Newton slope} $t_N(D,\phi_D,L_K):=\ord_p(\det\phi_D)$ and the \emph{Hodge slope} $t_H(D,\phi_D,L_K):=\dim_K L_K-\dim_K D_K$. Following Fontaine \cite{Fontaine79} and Rapoport-Zink~\cite[1.18]{RZ}, the point $L_K\in\CF$ is called \emph{weakly admissible} if
\begin{eqnarray*}
t_N(D,\phi_D,L_K)&=&t_H(D,\phi_D,L_K)\es=\es-d\qquad\text{and}\\
t_N(D',\phi_D|_{D'},L_K\cap D'_K)&\ge& t_H(D',\phi_D|_{D'},L_K\cap D'_K)
\end{eqnarray*}
for all $\phi_D$-stable $K_0$-subspaces $D'\subset D$. Let $\CF^\rig$ be the rigid analytic space, and $\CF^\an$ the $K_0$-analytic space in the sense of Berkovich~\cite{Berkovich,Berkovich2} associated with $\CF$. 

\begin{theorem} (Rapoport-Zink~\cite[Proposition 1.36]{RZ})\\
The set $\CF^\rig_{wa}:=\{\,L_K\in\CF^\rig:L_K \text{ is weakly admissible}\,\}$ is an open rigid analytic subspace. 
\end{theorem}

The space $\CF^\rig_{wa}$ is an example for the \emph{$p$-adic period domains} constructed more generally in \cite[Proposition 1.36]{RZ} for arbitrary filtered isocrystals. The proof of Rapoport and Zink even shows that $\CF^\rig_{wa}$ is the rigid analytic space associated to an open $K_0$-analytic subspace $\CF^\an_{wa}\subset\CF^\an$; see \cite[Proposition 1.3]{HartlRZ}. The period domain $\CF^\rig_{wa}$ contains the set of Grothendieck's problem~\ref{ProblemGroth}. To explain this let $\Nilp_W$ be the category of $W$-schemes on which $p$ is locally nilpotent. For an $S\in\Nilp_W$ we set $\bar S:=\Var(p)\subset S$.

\begin{theorem}\label{ThmRZ} (Rapoport-Zink~\cite[Theorem 2.16]{RZ})
The functor $\CG:\Nilp_W\to \Sets$
\begin{eqnarray*}
S&\longmapsto& \bigl\{\,\text{isomorphism classes of pairs }(X,\rho)\text{ where }X\text{ is a Barsotti-Tate}\\
&&\qquad\text{ group over $S$ and }\rho: \BX_{\bar S}\to X_{\bar S}\text{ is a quasi-isogeny}\,\bigr\}
\end{eqnarray*}
is pro-representable by a formal scheme locally formally of finite type over $W$.
\end{theorem}

To $\CG$ one can associate a rigid analytic space $\CG^\rig$ by Berthelot's construction~\cite{Berthelot96}, and also a $K_0$-analytic space $\CG^\an$. Rapoport and Zink~\cite[5.16]{RZ} construct a period morphism $\breve\pi_1^\rig:\CG^\rig\to\CF^\rig_{wa}$ as follows. By the theory of Grothendieck-Messing~\cite{Messing}, the universal Barsotti-Tate group $X$ over $\CG$ gives rise to an extension
\[
\xymatrix {
0\ar[r] & (\Lie X\dual)_{\CG^\rig}\dual \ar[r] & \BD(X_{\bar\CG})_{\CG^\rig}\ar[r] & \Lie X_{\CG^\rig}\ar[r] & 0
}
\]
of locally free sheaves on $\CG^\rig$. The quasi-isogeny $\rho:X_{\bar\CG}\to\BX_{\bar\CG}$ induces by the crystalline nature of $\BD(\,.\,)$ an isomorphism $\BD(\rho)_{\CG^\rig}:\BD(X_{\bar\CG})_{\CG^\rig}\isoto\BD(\BX)_{\CG^\rig}$ and the image $\BD(\rho)_{\CG^\rig}(\Lie X\dual)_{\CG^\rig}\dual$ defines a $\CG^\rig$-valued point of $\CF^\rig$. By \cite[5.27]{RZ} the induced morphism $\CG^\rig\to\CF^\rig$ factors through $\CF^\rig_{wa}$. This is the period morphism $\breve\pi_1^\rig$. The same construction also gives a $K_0$-analytic period morphism $\breve\pi_1^\an:\CG^\an\to\CF^\an$.

\begin{theorem}\label{ThmGrothQuest}(Rapoport-Zink, Colmez-Fontaine, Breuil, Kisin)\\
The set of Grothendieck's Problem~\ref{ProblemGroth} is contained in the subset $\CF^\rig_{wa}$. The latter also equals the image of the period morphism $\breve\pi_1^\rig:\CG^\rig\to\CF^\rig$ in the sense of sets.
\end{theorem}

\begin{proof}
Let $L_K=(\Lie X\dual)_K\dual$ be a point in Grothendieck's set, given by a Barsotti-Tate group $X$ over $\CO_K$ with $\BX\cong X_{\BF_p^\alg}$. Over $\CO_K/(p)$ this isomorphism lifts by rigidity \cite[Appendix]{Drinfeld2} to a quasi-isogeny $\rho:\BX_{\CO_K/(p)}\to X_{\CO_K/(p)}$ and $(X,\rho)$ gives a point of $\CG(\CO_K)$. By construction $\breve\pi_1^\rig(X,\rho)_K=L_K$. So the point $L_K$ belongs to the image of $\breve\pi_1^\rig$, which in turn lies in $\CF^\rig_{wa}$. It remains to show that every $K$-valued point $L_K\in\CF^\rig_{wa}$ for $K/K_0$ finite lies in the image of $\breve\pi_1^\rig$. By the theorem of Colmez-Fontaine~\cite{CF} the filtered isocrystal $(D,\phi_D,L_K)$ is admissible, i.e. arises from a crystalline $p$-adic Galois representation $V$. By Breuil~\cite[Theorem 1.4]{Breuil2} there is a Barsotti-Tate group $X$ over $\CO_K$ and an isomorphism $T_p X_K\otimes_{\BZ_p}\BQ_p\cong V$ if $p>2$. Kisin~\cite[Corollary 2.2.6]{Kisin06} extended Breuil's theorem to $p=2$ and reproved the Colmez-Fontaine Theorem. Fontaine's functor $D_\cris$ transforms the isomorphism $T_p X_K\otimes_{\BZ_p}\BQ_p\cong V$ into an isomorphism $\bar\rho_\ast:\bigl(\BD(X)_{K_0},\BD(\Frob_X)_{K_0},(\Lie X\dual)_K\dual\bigr)\isoto(D,\phi_D,L_K)$ of filtered isocrystals. This defines a quasi-isogeny $\bar\rho:\BX\to X_{\BF_p^\alg}$ which again by rigidity lifts to a quasi-isogeny $\rho:\BX_{\CO_K/(p)}\to X_{\CO_K/(p)}$. So $L_K$ lies in the image of the period morphism.
\end{proof}

Unfortunately an open rigid analytic subspace of $\CF^\rig$, like the image of $\breve\pi_1^\rig$, is not uniquely determined by its underlying set of points. This however is true for $K_0$-analytic spaces. Therefore it is natural to ask: What is the image of $\breve\pi_1^\an$? Examples of Genestier-Lafforgue and the author (see Example~\ref{Ex5.4} below) show that this image is in general smaller than $\CF^\an_{wa}$. As \cite[1.37 and 5.53]{RZ} indicate, Rapoport and Zink were aware of this phenomenon. We determine the image of $\breve\pi_1^\an$ in Section~\ref{Sect3}. But before we need to recall some of the rings of Fontaine Theory.

\section{Some of Fontaine's rings}

Let $K/K_0$ be a complete, rank one valued field extension, let $C$ be the completion of an algebraic closure of $K$, and let $\CO_C$ be its valuation ring. Starting out from $C$ various rings are defined in Fontaine Theory. For details on their construction see Colmez~\cite{Colmez}. We follow his notation.

We let $\wt\bE^+:=\bigl\{\,u=(u^{(n)})_{n\in\BN_0}:\es u^{(n)}\in\CO_C, (u^{(n+1)})^p=u^{(n)}\,\bigr\}$. 
With the 
multiplication $uv:=\bigl(u^{(n)}v^{(n)}\bigr)_{n\in\BN_0}$, the 
addition $u+v:=\bigl(\lim_{m\to\infty}(u^{(m+n)}+v^{(m+n)})^{p^m}\bigr)_{n\in\BN_0}$, and 
the valuation $v_\bE(u):=v_C(u^{(0)})$, $\wt \bE^+$ becomes a complete valuation ring of rank one with algebraically closed fraction field, called $\wt\bE$, of characteristic $p$.
Fix primitive $p^n$-th roots of unity $\epsilon^{(n)}$ such that $\epsilon:=(1,\epsilon^{(1)},\epsilon^{(2)},\ldots)$ is an element of $\wt\bE^+$. 

Let $\wt\bA:= W(\wt\bE)$
~and consider the automorphism $\phi=W(\Frob_p)$ of
$\wt\bA$. For an element $u\in\wt\bE$ let $[u]\in\wt\bA$ be its Teichm\"uller representative.

If $x=\sum_{i=0}^\infty p^i[x_i]\in\wt\bA$ then we set $w_k(x):=\min\{\,v_\bE(x_i):i\le k\,\}$. For $r>0$ let $\wt\bA^{(0,r]}\DS:=\bigl\{\,x\in\wt\bA:\es\lim_{k\to+\infty}w_k(x)+{\TS\frac{k}{r}}=+\infty\,\bigr\}$ and let $\wt\bB^{(0,r]}:=\wt\bA^{(0,r]}[\frac{1}{p}]$.

On $\wt\bB^{(0,r]}$ there is a valuation defined for $x=\sum_{i\gg-\infty}^\infty p^i[x_i]$ as\\
 $v^{(0,r]}(x)\es:=\es\min\{\,w_k(x)+\frac{k}{r}:\,k\in\BZ\,\}\;=\;\min\{\,v_\bE(x_i)+\frac{i}{r}:\,i\in\BZ\,\}$.

Let $\wt\bB^{]0,r]}$ be the Fr\'echet completion of $\wt\bB^{(0,r]}$ 
with respect to the family of semi-valuations $v^{[s,r]}(x):=\min\{\,v^{(0,s]}(x),v^{(0,r]}(x)\,\}$ for $0<s\le r$. The logarithm $t:=\log[\epsilon]$ converges to an element in $\wt\bB^{]0,1]}$.

Let $\wt\bB^\dagger_\rig:=\es\bigcup_{r>0}\wt\bB^{]0,r]}$. The homomorphism $\phi$ gives rise to a bicontinuous isomorphism $\phi:\wt\bB^{]0,r]}\isoto\wt\bB^{]0,r/p]}$ and thus induces an automorphism of $\wt\bB^\dagger_\rig$. It satisfies $\phi(t)=p\,t$.

Finally there is a homomorphism $\theta:\wt\bB^{(0,1]}\to C$ sending $\sum_{i\gg-\infty}^\infty p^i[x_i]$ to $\sum_{i\gg-\infty}^\infty p^ix_i^{(0)}$ which extends by continuity to $\wt\bB^{]0,1]}$. The element $t$ lies in the kernel of $\theta$.

\bigskip

\begin{definition}
A \emph{$\phi$-module over $\wt\bB^\dagger_\rig$} is a finite free $\wt\bB^\dagger_\rig$-module $\CM$ together with a $\phi$-linear automorphism $\phi_\CM:\CM\to\CM$. 
\end{definition}

The following structure theorem was proved by Kedlaya~\cite[Theorem 4.5.7]{Kedlaya}.

\begin{theorem} \label{ThmKedlaya}
Every $\phi$-module $\CM$ over $\wt\bB^\dagger_\rig$ is isomorphic to $\bigoplus_i\CM_{c_i,d_i}$ where $\CM_{c,d}=(\wt\bB^\dagger_\rig)^{\oplus d}$, $\phi_{\CM_{c,d}}=
\left(\raisebox{0.55cm}{\xymatrix @R=0pc @C=0.4pc{0 \ar@{.}[ddrr]& & p^c \\ 1 \ar@{.}[dr] \\ & 1 & 0 }}\right)\cdot\phi$ \mbox{for $c,d\in\BZ$ with $(c,d)=1$ and $d>0$.}
\end{theorem}

One defines the \emph{degree of $\CM$} as $\deg\CM:=\sum_i c_i$.

\section{The construction of $\CF^\an_a$} \label{Sect3}

As above let $D=\BD(\BX)_{K_0}$ and $\phi_D=\BD(\Frob_\BX)_{K_0}$.
Let $(\calD^{]0,1]},\phi_\calD):= (D,\phi_D)\otimes_{K_0}\wt\bB^{]0,1]}$ and consider the morphism $1\otimes\theta:\calD^{]0,1]}\to D\otimes_{K_0} C$. By a variant of a construction of Berger~\cite[\S II]{Berger04} every point $L=L_K\in\CF^\an$, with values in a field $K$ as in Section 2, defines a $\phi$-module over $\wt\bB^\dagger_\rig$ as follows.

\begin{proposition} \label{Prop3.1} (\cite[Proposition 4.1]{HartlRZ})\\
There exists a uniquely determined $\wt\bB^{]0,1]}$-submodule $t\,\calD^{]0,1]}\subset\CM_L^{]0,1]}\subset\calD^{]0,1]}$ such that $(1\otimes\theta)\bigl(\CM_L^{]0,1]}\bigr)=L_K\otimes_K C$ and $\phi_\calD:\CM_L^{]0,1]}\isoto\CM_L^{]0,1]}\otimes_{\wt\bB^{]0,1]}}\wt\bB^{]0,1/p]}$ is an isomorphism. In particular $\CM_L^{]0,1]}$ defines a $\phi$-module $\CM_L:=\CM_L^{]0,1]}\otimes_{\wt\bB^{]0,1]}}\wt\bB^\dagger_\rig$ over $\wt\bB^\dagger_\rig$.
\end{proposition}

The following results are proved in \cite{HartlRZ}.

\begin{theorem} (\cite[Theorem 4.4]{HartlRZ})\\
$\deg\CM_L=t_N(D,\phi_D,L_K)-t_H(D,\phi_D,L_K)$.
\end{theorem}

Consider the subset $\CF^\an_a:=\{\,L\in\CF^\an:\CM_L\cong\CM_{0,1}^{\oplus h}\,\}$ of the $K_0$-analytic space $\CF^\an$.

\begin{theorem} \label{Thm3.3} (\cite[Theorem 5.2]{HartlRZ})\\
The set $\CF^\an_a$ is an open $K_0$-analytic subspace of $\CF^\an_{wa}$.
\end{theorem}

\begin{remark} (on the proof of Theorem~\ref{Thm3.3}. See also Remark~\ref{Rem3.8} below.)\\
The inclusion $\CF^\an_a\subset\CF^\an_{wa}$ is seen as follows. If $D'\subset D$ is a $\phi_D$-stable $K_0$-subspace then $(D',\phi_D|_{D'},L_K\cap D_K)$ defines by Proposition~\ref{Prop3.1} a $\phi$-submodule $\CM_L'\subset\CM_L$. Since $\CM_{0,1}^{\oplus h}$ is ``semistable'' of slope zero we conclude by \cite[Lemma 3.4.8]{Kedlaya} that
\[
t_N(D',\phi_D|_{D'},L_K\cap D'_K)\;-\; t_H(D',\phi_D|_{D'},L_K\cap D'_K)\es=\es\deg\CM_L'\es\ge\es\deg\CM_L\es=\es0
\]
with equality if $D'=D$. Hence $\CF^\an_a\subset\CF^\an_{wa}$. On the other hand Berger's proof~\cite{Berger04} of the Colmez-Fontaine Theorem shows that this inclusion induces a bijection on rigid analytic points (with $K/K_0$ finite), or more generally points of $\CF^\an_{wa}$ with with values in a discretely valued field $K$.
\end{remark}

\begin{theorem} 
$\CF^\an_a$ is the image of the period morphism $\breve\pi_1^\an:\CG^\an\to\CF^\an_{wa}$.
\end{theorem}

\begin{proof} Let $L_K\in\CF^\an$ be a $K$-valued point in the image of $\breve\pi_1^\an$ with $K/K_0$ not necessarily finite. So $L_K=\BD(\rho)_K(\Lie X\dual)_K\dual$ for a Barsotti-Tate group $X$ over $\CO_K$ and a quasi-isogeny $\rho:\BX_{\CO_K/(p)}\to X_{\CO_K/(p)}$. Then the Tate module $T_pX_K$ of $X$ induces an injection $\CM_{0,1}^{\oplus h}\cong T_pX_K\otimes_{\BZ_p}\wt\bB^\dagger_\rig\hookrightarrow\CM_L$ which must be an isomorphism by reasons of degree. This was proved in \cite[Proposition 6.1]{HartlRZ}. Thus $L_K\in\CF^\an_a$. 

Conversely let $L_K\in\CF^\an_a$. Then the morphism $(\wt\bB^\dagger_\rig)^{\oplus h}\isoto\CM_{L_K}\hookrightarrow D\otimes_{K_0}\wt\bB^\dagger_\rig$ is represented, with respect to a $K_0$-basis of $D$, by a matrix $M\in M_h(\wt\bB^\dagger_\rig)$ with $tM^{-1}\in M_h(\wt\bB^\dagger_\rig)$. Then in fact $M,tM^{-1}\in M_h(\wt\bB^+_\rig)\subset M_h(\bB^+_\cris)$ by \cite[Proposition I.4.1]{Berger04b}. For notation see \cite[\S I.1]{Berger04b} or \cite[\S2]{HartlRZ}. So $M$ defines an isomorphism $\bB_\cris^{\oplus h}\isoto D\otimes_{K_0}\bB_\cris$ compatible with Frobenius, which maps $(\bB_\cris^+)^{\oplus h}$ onto the preimage of $L_K\otimes_K C$ under the map $\id\otimes\theta:D\otimes_{K_0}\bB_\cris^+\to D\otimes_{K_0}C$. This means that $(D,\phi_D,L_K)\dual$ is admissible in the sense of Faltings' \cite[Definition 1]{Faltings07}. Note that Faltings uses
 contravariant Dieudonn\'e modules. By \cite[Theorems 9 and 14]{Faltings07}, $L_K=\BD(\rho)_K(\Lie X\dual)_K\dual$ for a Barsotti-Tate group $X$ over $\CO_K$ and a quasi-isogeny $\rho:\BX_{\CO_K/(p)}\to X_{\CO_K/(p)}$, hence $L_K$ lies in the image of $\breve\pi_1^\an$.
\end{proof}

There are many examples showing that only in rare cases $\CF^\an_a=\CF^\an_{wa}$. We mention one here. Similar examples are due to A.\ Genestier and V.\ Lafforgue.

\begin{bigexample} \label{Ex5.4}
Let $D=K_0^{\oplus 5}$ and $\phi_D=\left(\raisebox{0.55cm}{\xymatrix @R=0pc @C=0.4pc{0 \ar@{.}[ddrr]& & p^{-3} \\ 1 \ar@{.}[dr] \\ & 1 & 0 }}\right)\!\cdot\phi$\,. Then $h=5\,,\, d=3$ and $\CF=\Grass(2,5)$. Since the isocrystal $(D,\phi_D)$ is simple $\CF^\an_{wa}=\CF^\an$. Let $L=L_K\in\CF^\an$. Then
\[
\CM_{-3,5}\es=\es \calD\es\supset\es\CM_L\es\supset\es t\,\calD\es\cong\es \CM_{2,5}\,.
\]
By Theorem~\ref{ThmKedlaya}, $\CM_L\cong\bigoplus_i \CM_{c_i,d_i}$ with $\sum_i d_i=\rk\CM_L=5$ and $\sum_i c_i=\deg\CM_L=0$. Moreover by \cite[Lemma 3.4.8]{Kedlaya} all the weights ${c_i}/{d_i}$ must lie between $-3/5$ and $2/5$. So either $\CM_L\cong\CM_{0,1}^{\oplus 5}$ or $\CM_L\cong\CM_{-1,2}\oplus\CM_{1,3}$. 

Now one easily checks that $\Hom_\phi\bigl(\CM_{-1,2},\calD\bigr)\;=\;\Hom_\phi\bigl(\CM_{-1,2},\CM_{-3,5}\bigr)\;=$
\[
\Biggl\{\,A\;=\;\left(  \begin{array}{cc}
\phi^5(x) & x \\ \phi^{11}(x) & \phi^6(x) \\ \phi^{17}(x) & \phi^{12}(x) \\ \phi^{23}(x) & \phi^{18}(x) \\ \phi^{29}(x) & \phi^{24}(x)   \end{array}\right)
\DS:\es x\;=\;\sum_{\nu\in\BZ}p^\nu\phi^{-10\nu}([u]),\;
u\in\bE,\;0< v_\bE(u)<\infty
\Biggr\}\,.
\]
The bad situation $\CM_L\cong\CM_{-1,2}\oplus\CM_{1,3}$ occurs if and only if $L_K$ is generated by the columns of such a matrix $\theta(A)\in C^{5\times 2}$, since then the homomorphism $A$ factors through $\CM_L$ and this forbids $\CM_L\cong\CM_{0,1}^{\oplus 5}$ by \cite[Lemma 3.4.8]{Kedlaya}.
Since obviously such $L_K$ exist, this proves that the inclusion $\CF^\an_a\subset\CF^\an_{wa}$ is strict.
\end{bigexample}

\forget{
\begin{remark} 
The example can easily be adapted to the general situation showing that only in rare cases $\CF^\an_a$ will equal $\CF^\an_{wa}$. Moreover, the question whether the two sets are equal depends largely on the combinatorics of the weights.
\end{remark}
}

\begin{remark} \label{Rem3.8} (on the proof of Theorem~\ref{Thm3.3}.)
One can explain the idea for the proof of Theorem~\ref{Thm3.3} by means of this example. Namely the complement $\CF^\an\setminus\CF^\an_a$ is the image of the continuous map from the compact set $\{u\in\wt\bE^+:\;1\le v_\bE(u)\le p^{10}\,\}$ given by $u\mapsto\theta(A)$.
\end{remark}


\textit{Acknowledgements.} The author is grateful to 
L.\ Berger, 
J.-M.\ Fontaine, V.\ Berkovich, A.\ Genestier, M.\ Kisin, V.\ Lafforgue, and M.\ Rapoport for various helpful discussions and their interest in this work.

\vfill

\noindent
Urs Hartl\\
University of M\"unster\\
Institute of Mathematics\\
Einsteinstr.~62\\
D -- 48149 M\"unster\\
Germany\\[1mm]
E-mail: 
uhartl@uni-muenster.de

\end{document}